OpenStudy (anonymous):

what is 3/18 in the simplest form

OpenStudy (mathlegend):

What is the highest number that can go into 3 and 18? You can start listing factors of each if that helps.

OpenStudy (mathlegend):

For instance... 2/8 Factors of 2: 1*2 Factors of 8: 1*8 2*4 What is the greatest number that we have in common here?

OpenStudy (anonymous):

3/18 3 = 3 * 1 18 = 3 * 6 (3 / 3) * (1 / 6) 3 / 3 = 1 1/6

OpenStudy (anonymous):

1/6

OpenStudy (mathlegend):

You should be trying to guide users to answers. Assuring the individual understands what you are doing. Justifying why you do what you do. @britneyspears You should not just be giving direct answers @SiennaSurfer1 please read the code of conduct. :)

OpenStudy (the_fizicx99):

You can do that but, you can look at that biggest number goes into the smallest number. For example: 14/28 The highest number that goes into 14 is and 7, also look at the other number, see if its also a factor of that number as well. 7 is a factor of 28! So 14 becomes 2 and 28 becomes 4 2/4 <-- this can also be further simplified, by 2. 2 goes into 2 1 time and 2 goes into 4 2 times. 1/2.

OpenStudy (the_fizicx99):

>.> ^

OpenStudy (anonymous):

Someone already worked it out. I was just justifying their answer.

OpenStudy (the_fizicx99):

My own arrow of shame ggs